### v1

# Benchmark the Linguistic Competence of Language Models

Abstract
We introduce Holmes, a benchmark to assess the linguistic competence of language models (LMs) – their ability to grasp linguistic phenomena. Unlike prior prompting-based evaluations, Holmes assesses the linguistic competence of LMs via their internal representations using classifier-based probing. In doing so, we disentangle specific phenomena (e.g., part-of-speech of words) from other cognitive abilities, like following textual instructions, and meet recent calls to assess LMs’ linguistic competence in isolation. Composing Holmes, we review over 250 probing studies and feature more than 200 datasets to assess syntax, morphology, semantics, reasoning, and discourse phenomena. Analyzing over 50 LMs reveals that, aligned with known trends, their linguistic competence correlates with model size. However, surprisingly, model architecture and instruction tuning also significantly influence performance, particularly in morphology and syntax. Finally, we propose FlashHolmes, a streamlined version of Holmes designed to lower the high computation load while maintaining high-ranking precision.
1 Introduction
Linguistic competence is the unconscious understanding of language, like grasping grammatical rules (Chomsky, 1965). As language models (LMs) are trained on simple tasks like next word prediction (Brown et al., 2020), one might naturally wonder: What is the linguistic competence of LMs, and how do they differ? To answer such questions, benchmarks estimate cognitive abilities by providing textual instructions and evaluate LMs’ responses, as done for mathematical reasoning (Cobbe et al., 2021) or factual knowledge (Petroni et al., 2019, 2020). However, they conflate latent abilities (like following provided instructions) with those under test, such as understanding specific linguistic phenomena, e.g., syntactic structures (Liang et al., 2023). As this entanglement makes it infeasible to draw definitive conclusions about distinct abilities (Hu and Levy, 2023), recent studies call to assess the linguistic competence of LMs comprehensively and in isolation (Lu et al., 2023; Mahowald et al., 2024).
Figure 1: A subset of Holmes rankings ($\downarrow$) for various evaluated LMs. FLAN-UL2 outperforms the others overall, while different LMs prevail for the five distinct types of linguistic phenomena.
Figure 2: Overview of Holmes (left) with the five phenomena types (right) and an example of probing-based evaluations for part-of-speech: encoding the input tokens and predicting the POS tag for cucumber, here NN.
In this work, we introduce the Holmes (Figure 2). A benchmark to assess the linguistic competence of LMs (Figure 1) regarding numerous linguistic phenomena. To fully disentangle the understanding of these phenomena and other abilities of LMs, we use classifier-based probing (Tenney et al., 2019a; Hewitt and Manning, 2019; Belinkov, 2022). A method that uses the LMs’ internal representations of text inputs to train linear models (probes) to predict specific aspects of phenomena, such as words’ part-of-speech (POS). We then approximate the LMs’ grasp of these phenomena using the probes’ performance, rigorously verified using control tasks (Hewitt and Liang, 2019) and from an information theory perspective (Voita and Titov, 2020). With this particular and comprehensive scope, we thoroughly address the initially raised questions as follows:
Meta-Study (§ 3)
The review of over 270 probing studies reveals a gap in comprehensively evaluating linguistic competence. Despite covering over 200 probing tasks and 150 LMs, individual studies focus on particular tasks and LMs. As a result, only three LMs were probed on over 20% of the tasks, and one single task was evaluated for more than 20% of the reviewed LMs. Notably, recent large LMs are significantly underrepresented.
Benchmark (§ 4)
To address this identified deficiency, Holmes offers a structured framework to assess the English linguistic competence of LMs comprehensively. It features 208 distinct datasets covering morphology, syntax, semantics, reasoning, and discourse phenomena, including previously underrepresented ones like negation or rhetoric in text (Liang et al., 2023).
Results and Analysis (§ 5)
From assessing 59 LMs (Figure 1), we find that no single one consistently excels the others and that their linguistic competence is more pronounced for morphology and syntax than the other phenomena types. Instead, we find model size, model architecture, and instruction tuning fundamentally affect their linguistic competence.
First, LMs’ linguistic competence, particularly for morphology and syntax, scales with their model size. This generalizes previous findings (Tenney et al., 2019b; Zhang et al., 2021) beyond LMs with 350 million parameters. Second, contrary to prompting evaluations (Lu et al., 2023) and aligned with other work (Waldis et al., 2024a; Gautam et al., 2024), model architecture is critical. The linguistic competence of decoder-only LMs is less pronounced, and even 70 billion does not allow them to encode linguistic phenomena of words with comparable strength to encoder-only LMs of a similar size. Third, while previous studies focused on aligning LMs with human interactions through instruction tuning (Ouyang et al., 2022; Touvron et al., 2023; Zhou et al., 2023), we show for the first time its effect on their linguistic competence. It improves morphology and syntax but has mixed effects for the other types of phenomena. Lastly, we contrast the results of Holmes with OpenLLM (Beeching et al., 2023), an extensive LM benchmark focusing on user-centered applications like mathematical reasoning. We find that Holmes provides a unique but supplementary perspective, as rankings partly align, especially for reasoning-related phenomena.
Efficiency (§ 6)
Finally, to mitigate the heavy computational burden of evaluating a new LM on Holmes, we form the streamlined version FlashHolmes by selectively excluding samples not significantly influencing overall rankings (Perlitz et al., 2023). Specifically, FlashHolmes approximates Holmes rankings with high precision while requiring only ~3% of the computation.
We summarize our contributions as follows:
  • Benchmark. Holmes comprehensively and thoroughly assesses the linguistic competence of LMs in isolation, providing substantial ground for advancements in NLP.
  • Empirical insights. Extensive experiments reveal that LMs’ linguistic competence is more pronounced for morphology and syntax, and size, architecture, and instruction tuning are crucial for LM differences.
  • Ease of use. We provide tools to interactively explore Holmes results and straightforward code to evaluate upcoming LMs with efficiency in mind (FlashHolmes).
2 Preliminaries
Language Models (LMs)
Language Models compute probabilities for word sequences $i$, enabling tasks such as classifying $i$, textual comparisons between $i$ and another sequence $i^{\prime}$, and text generation based on $i$. We consider LMs as any model producing representations of input $i$, regardless of their specific type: sparse like bag-of-words (Harris, 1954); static such as GloVe (Pennington et al., 2014); or contextualized transformer-based LMs (Devlin et al., 2019; Raffel et al., 2020).
Linguistic Competence
Following Chomsky (1965), linguistic competence is defined as the unconscious knowledge of language, encompassing the understanding of specific linguistic phenomena, including word dependencies and their distinct parts of speech (POS).
Linguistic Phenomena
We define the linguistic competence of LMs as their ability to understand a diversity of linguistic phenomena. Specifically, we focus on five phenomena types: morphology, the structure of words; syntax, the structure of sentences; semantics, the meaning of words; reasoning, the use of words in logical deduction and other related phenomena like negation or speculation; discourse, the context in text like rhetorical structure. Following Mahowald et al. (2024), we categorize these phenomena types into two groups: morphology and syntax are formal phenomena, which include understanding grammatical rules and statistical patterns, while functional ones (semantics, reasoning, and discourse) focus on practical abilities like interpreting text sentiment or detecting the existence of speculation.
Datasets
We define a dataset as text examples and labels covering a specific aspect of a linguistic phenomenon, like words and their POS tag. Typically, these labels are highly unambiguous to assess the specific aspect under test in isolation.
Probes
Using probes, we empirically assess the linguistic competence of LMs regarding the featured linguistic phenomena in Holmes. To this end, we employ probing tasks using the widely recognized classifier-based probing method (Tenney et al., 2019a; Hewitt and Manning, 2019; Belinkov, 2022), or known as diagnostic classifiers (Veldhoen et al., 2016; Giulianelli et al., 2018). Running such a probing task involves training a probe (linear model) using the specific dataset to test a distinct aspect of a linguistic phenomenon in isolation. Therefore, we feed the text examples, encoded with a given LM, as training inputs. Subsequently, we use the probe’s performance to approximate how an LM understands the specific linguistic phenomenon under test. With a higher score, we assume the embeddings embody patterns relevant to this phenomenon, which enhances the accuracy (Tenney et al., 2019b).
3 Meta-Study
In this section, we survey 274 studies (§ 3.1), probing LMs’ linguistic competence. We analyze these studies regarding their evolution, covered probing tasks and LMs (§ 3.2), and identify the apparent need for consolidating existing resources (§ 3.3).
3.1 Scope
We analyze 28k papers ($P$) from 2015 to August 2023 of major NLP conferences (TACL, ACL, AACL, COLING, EACL, EMNLP, NAACL, and corresponding workshops) expanded with selected work from other venues such as ICLR. To identify relevant work, we follow a semiautomatic approach. First, we automatically select papers based on their meta-data and full text. [^1] We use PyPDF2 v3.0.0, DBLP and semanticscholar API. We select a total of 493 candidate papers matching at least one of the following three criteria ($P^{\prime}=\{\forall p\in P|p\in P_{1}\cup p\in P_{2}\ \cup p\in P_{3}\}$):
$\mathbf{P_{1}}$: papers contain probing or probe in the title.
$\mathbf{P_{2}}$: papers contain probing or probe in the abstract and at least five times in the main content.
$\mathbf{P_{3}}$: papers contain probing or probe at least ten times in the main content.
We manually verified these automatically curated candidates ($P^{\prime}$) and found 274 relevant papers ($P_{r}$).We selected them as they either evaluate LMs regarding one or more linguistic phenomena as part of the analysis or as a main contribution. This involves filtering papers using the term probing in other senses, such as probing hash tables (Bogoychev and Lopez, 2016).
3.2 Analysis
Next, we analyze these 274 relevant studies ($P_{r}$).
i) Scattered evolution calls for consolidation.
First, we analyze the evolution of the relevant studies. Figure 3 relates how these studies cite each other (probing citations $C_{p}$) compared to other gathered citations (general citations $C_{g}$). Colorized, we show the ratio $\alpha$ between these two measures $\alpha=\frac{|C_{p}|+1}{|C_{g}|+1}$. First, only a fraction of the works gained general attention, as 16 papers exceeded 200 general citations. Further, probing works cite each other rather sparsely, with an average probing citation ratio of $\alpha=0.1$. Therefore, we see other fields are paying little attention to the linguistic competence of LMs. Paired with scattered citation patterns, we identify the need to consolidate existing resources to solidly ground research in this field.
Figure 3: Citation analysis considering probing citations originating from the set of relevant work and every other citation (general citations). The color scale indicates the ratio ($\alpha$) between them.
ii) Probing work prioritizes tasks and analytics over methods.
We categorize the selected work according to their probing focus: methodological, new methods, like control tasks (Hewitt and Liang, 2019) or minimum description length (Voita and Titov, 2020); task-focused assessing specific linguistic phenomena as main contributions, such as discourse relations in text (Koto et al., 2021); and analytical using probing tasks to analyze LMs, such as the impact of pre-training data (Zhang et al., 2021). Figure 4 shows: the majority (51.8%) of studies focus on specific probing tasks like numeric scales (Zhang et al., 2020), or morphosyntactic (Shapiro et al., 2021); 35.7% use probing as a supplementary analytical tool, for example, analyzing the effect of fine-tuning (Mosbach et al., 2020a; Zhu et al., 2022a); 12.5% address methodological problems related to probing (Wu et al., 2020; Immer et al., 2022; Zhu et al., 2022b).
Figure 4: Categorization of the selected studies by their focus and their conducted probing method.
iii) The dominance of classifier-based probing.
Next, we analyze the specific employed probing method: classifier, using linear or shallow models to probe internal representations of LMs, as demonstrated in Tenney et al. (2019a); mask, letting LMs fill gaps to verify linguistic phenomena, as shown in Talmor et al. (2020) or Warstadt et al. (2020); attention, which relies on attention patterns, as used in Pandit and Hou (2021) for bridging; and other, methods not belonging to the previous three categories, such as dimension selection (Torroba Hennigen et al., 2020). Most studies utilize the classifier-based probing method (74%), 20% conduct mask-based probing, and only a minority of work ($\sim$ 3%) considers attention patterns or other approaches.
iv) Tasks and LMs are barely broadly evaluated.
Finally, we analyze which tasks and LMs the relevant probing studies consider. For example, Tenney et al. (2019b) considers BERT and probes POS tagging, semantic-role labeling (SRL), and other ones. Aggregated over all studies, we found a broad coverage of 289 unique tasks and 161 distinct LMs. Below, we delve into the details and highlight noteworthy findings.
We analyze how LMs and tasks are considered jointly in Figure 5. Despite the broad coverage, single studies, including fundamental ones, maintain a particular focus and consider only a fraction of LMs and tasks. For example, while most tasks (72%) were assessed on BERT, RoBERTa’s coverage has already declined to 42%. Conversely, part-of-speech tagging (POS), the most probed task, was only evaluated on 23% of the LMs, for example, not covering prominent examples like BART Lewis et al. (2020). Notably, more recently released larger and powerful LMs, like PYTHIA (Biderman et al., 2023), UL2 (Tay et al., 2023), or LLAMA-2 (Touvron et al., 2023), and instruction-tuned LMs (FLAN-T5 (Chung et al., 2022), LLAMA-2-Chat (Touvron et al., 2023), or TK-Instruct (Wang et al., 2022) are missing almost entirely, with single more recent exceptions (Hu and Levy, 2023; Waldis et al., 2024a). Again, these insights underscore the need to consolidate existing resources for more dense coverage. This is further evident when considering Figure 5, where we sort LMs and tasks according to how often they were mentioned in the relevant works. Then, we plot their cumulative coverage concerning all mentions. For example, considering all studies (red line), the top-10 most mentioned LMs account for 80% of all LMs mentions (black dot). In contrast, the other 151 unique LMs account for only 40%. Comparing the paper focus, we see that methodological studies rely only on 24 LMs and 36 tasks. In contrast, task-focused and analytical work covers a similar number of LMs (91 and 99, respectively). However, due to their distinct focus, task-focused studies cover significantly more tasks (202) than analytical ones (115).
Figure 5: Overview of how many tasks single LMs cover and vice versa - single examples are highlighted.
Figure 6: Cumulative coverage of LMs and tasks, considering all relevant studies and their focus.
3.3 Summary
This meta-study emphasizes the need to consolidate existing resources for a comprehensive assessment of the linguistic competence of LMs — a manifold but rather blind spot in evaluation research. Apart from more thorough evaluations, such a stimulus can significantly boost future research, as happened in computer vision with ImageNet (Deng et al., 2009) or in NLP with GLUE and SuperGLUE (Wang et al., 2019a, b).
4 Holmes Benchmark
With Holmes, we provide an extensive ground to tackle these identified deficiencies in the existing literature and comprehensively investigate the English linguistic competence of LMs. Specifically, Holmes features 208 datasets addressing distinct aspects of 66 phenomena covering morphology, syntax, semantic, reasoning, and discourse.
4.1 Datasets
To feature a total of 208 unique datasets, we leverage existing and established resources like OntoNotes (Weischedel et al., 2013), English Web Treebank (Silveira et al., 2014), or BLIMP (Warstadt et al., 2020) and create datasets addressing phenomena like the POS of words, their dependencies or determine the linguistic acceptability of sentences. Further, we include a range of less employed data, addressing contextualization of words (Klafka and Ettinger, 2020), reasoning (Talmor et al., 2020), semantic decomposition (White et al., 2016; Rudinger et al., 2018a, b; Govindarajan et al., 2019; Vashishtha et al., 2019), grammatical knowledge (Huebner et al., 2021), bridging (Pandit and Hou, 2021), and rhetorical (Carlson et al., 2001) and discourse (Webber et al., 2019) structure in text. Finally, we cover rarely probed phenomena like negation (Szarvas et al., 2008; Konstantinova et al., 2012; Vahtola et al., 2022), or word complexity (Paetzold and Specia, 2016).
4.2 Structure
Apart from the comprehensive scope, Holmes provides a clear structure for specific evaluations on different levels of aggregation. We first group the datasets according to the linguistic phenomena addressed. Then, we categorize these phenomena into their previously introduced type (see § 2) - morphology, syntax, semantics, reasoning and discourse. We rely on the categorization provided by the specific studies whenever given. The detailed categorization is given in § A.3.
4.3 Experimental Setup
Holmes evaluation follows the primarily used classifier-based probing paradigm, as described in § 2. Considering the internal representations allows us to maximally disentangle the understanding of distinct linguistic phenomena from each other and from other cognitive abilities (like following textual instructions). Further, this method allows us to assess any type of LMs, including sparse, static, or contextualized ones. Based on the specific dataset, we either select the embeddings of the specific input tokens (like single words for POS tagging) or average embeddings across a span or the whole sentence. We define a probing task as training a probe $f_{p}$ (linear model without intermediate layers) using these embeddings as inputs and the dataset labels as training signals. If not defined in the original data, we divide the dataset samples into train/dev/test split following a ratio of 70/10/20. We repeat this procedure five times using different random seeds and aggregate the results afterward.
4.4 Evaluations
We approximate how well an LM encodes specific linguistic phenomena using the absolute prediction performance of the probes. In addition, we rigorously evaluate the reliability of probing results using control tasks and from an information theory perspective (Voita and Titov, 2020; Hewitt and Liang, 2019). Different from commonly used prompting assessments, this particular evaluation protocol refrains from known fallacies in which the results and conclusions are sensible with specific instructions (Mizrahi et al., 2024; Min et al., 2022) or few-shot examples (Lu et al., 2023).
Task Score Metric
Based on a dataset’s specific task type, we use a corresponding performance measure, macro $F_{1}$ for classification or Pearson correlation for regression. In addition, we calculate the standard deviation $\sigma$ of the probe across multiple seeds. A lower $\sigma$ indicates a better encoding of a given linguistic phenomenon since the measurement is robust to noise. Further, we use the task score for ranking-based evaluation of all evaluated LMs $L=\{l_{1},...,l_{m}\}$ within Holmes. We calculate the mean winning rate $mwr$ (in percentage), telling us how many times one LM $l_{1}$ wins against others (Liang et al., 2023). With a higher $mwr$, we assume an LM encodes tested linguistic phenomena better than others.
Compression
Next, we evaluate the probes’ reliability from an information-theoretic perspective. Following Voita and Titov (2020), we use the compression $I$ to measure how well a probe compresses input data. A higher $I$ means fewer bits are needed, indicating that the given linguistic phenomenon is more clearly encoded in the embeddings.
Selectivity
A reliable probe should grasp patterns relevant to the tested phenomena in the internal representations of LMs but should not be able to learn anything else. Therefore, we expect high performance when evaluating the specific dataset but low performance when we randomize training signals. We check this using control tasks introduced in Hewitt and Liang (2019). Specifically, we calculate the selectivity $S$ as the difference between the probe trained with the original labels $y$ and the control task where we train the probe with randomly assigned labels $y^{\prime}$. With a higher $S$, we assume the detected patterns are relevant for the specific phenomena under test, as random patterns do not lead to similar performance.
5 Holmes Results
Using Holmes, we evaluate a diverse collection of 59 LMs. [^2] Find a complete list in Appendix § A.2. Using the results of these extensive experiments, we first answer the research question: what is the linguistic competence of LMs? In doing so, we discuss the reliability of results (i), the linguistic competence of LMs concerning the unique structure of Holmes (ii), and how these results relate to other downstream abilities (iii). Subsequently, we examine how linguistic competence varies among LMs, as we find LMs prevailing for different types of linguistic phenomena (Figure 1) and delve into the effects of model architecture (iv), size (v), and instruction tuning (vi).
i) The reliability of Holmes.
First, we show the reliability of probing-based evaluation, using deviation $\sigma$, compression $I$, and selectivity $S$ results in Figure 7. Single outliers are datasets that are too hard for all LMs, as the sample size is too small, or the linguistic phenomena under test are too complex, as the ability to detect spans causes speculations in a text. We average these metrics for every dataset across all LMs. Note, for selectivity, we consider only base-sized model (10m-200m parameters) for computational efficiency.
First, we found a low average deviation ($\sigma=0.02$), indicating the high reliability of probes across random seeds. These results also highlight the stability of probing results, compared to prompting-based ones where results across many paraphrased prompts lead to a deviation of $\sigma=0.07$ reported in Mizrahi et al. (2024). Next, substantial compression (average $I=1.9$) and selectivity (average $S=0.31$) further confirm the probes’ reliability. Interestingly, one identifies two parallel trends for selectivity. Harder datasets with many labels, like POS tagging, are arranged around a selectivity of 0.1 to 0.4 and a task metric of 0.3. In contrast, for easier binary classification tasks (such as linguistic applicability), we observe selectivity around 0.2 to 0.5 and a task metric of 0.6 to 0.9. Further, we measure a significant ($p<0.05$) positive correlation between the task metrics and the compression ($\tau=0.64$) and selectivity ($\tau=0.65$). This further confirms our reliability assumption and allows us to trust the task metric as the primary evaluation measure.
Figure 7: Reliability evaluation using deviation, compression (log), and selectivity on the y-axis for all 208 probing datasets. The x-axis represents the task metrics (either person correlation or macro $F_{1}$).
ii) The story of Holmes.
We focus on what Holmes tells us in general and regarding formal and functional phenomena, as defined in § 2. We report in Figure 8 the task metric, discriminability, and selectivity, averaged for every phenomena type. Note, discriminability (Rodriguez et al., 2021) quantifies the alignment of LMs ranking of one specific dataset compared to the overall rankings using the Kendall Tau correlation. Considering these three metrics, all tested LMs strongly encode formal phenomena (morphology and syntax), which often depend on the local neighborhood of words. Therefore, we assume that LMs approximate these co-occurrences during pre-training with high precision. For example, the specific POS tag of a word, like man (noun), primarily depends on its surroundings, such as the frequent predecessor the. In contrast, LMs encode less information about functional phenomena (semantics, reasoning, and discourse) since they show a relatively low performance regarding the task metric. For these functional phenomena, we assume more complex co-occurrences are required to capture the broad context in language, such as the rhetorical relation of two distant text spans. Despite these differences between formal and functional phenomena types, they contribute to the benchmark in a balanced way. A low to medium discriminability indicates that none of these types of linguistic phenomena dominates the overall LM rankings.
This balanced influence of the five phenomena types is further visible when considering their ranking correlations (Figure 9, left). A high average correlation of $67.8\pm 6.6$ with the overall results (last column/row) hints that they are facets of a broader occurrence but share common characteristics. Still, breaking into categories is meaningful, as the phenomena types (first five columns/rows) are medium correlated (average of $53.9\pm 14.5$). Analyzing the results of phenomena types further highlights the value of this distinction. While results of morphology and syntax are similarly correlated with the overall results ($68.2$ and $70.2$), their direct correlation ($69.1$) indicates their supplementary nature. Further, discourse results show the lowest correlation with others ($44.8\pm 16.1$), indicating the particular scope.
Figure 8: Average task metric, difficulty, and discriminability for each phenomena type. The dashed lines show the average measure over all datasets.
iii) The companions of Holmes.
We analyze how the results of Holmes and those from other evaluations focusing on downstream applications align (Figure 9, right). We select the OpenLLM benchmark (Beeching et al., 2023), as it covers a wide range of open LMs, in contrast to others like HELM (Liang et al., 2023). First, Holmes and OpenLLM results of jointly evaluated LMs are medium correlated, hinting that the linguistic competence of LMs is partly aligned with their downstream abilities. The nature of this alignment is further evident when focusing on morphology, reasoning, and discourse. Interestingly, and in contrast to syntax and semantics, their correlation to the OpenLLM and Holmes overall results is similar. Therefore, these three phenomena presumably represent skills that are more tested in the general benchmarks. These correlation patterns are consistent across the three most meaningful OpenLLM datasets (MMLU, TruthfulQA, and GSM8K). As TruthfulQA shows lower correlations with the linguistic phenomena and other datasets within OpenLLM, we presume this dataset captures distinctly different skills (possibly knowledge). [^3] Further, it’s also known that we need to expect this dataset to be fully leaked (Balloccu et al., 2024). These insights show how different benchmarks provide a different scope and supplement themselves simultaneously. Further, the above analysis shows, again, the value of assessing the linguistic competence of LMs across different phenomena types, for fine-grained analyses.
Figure 9: Kendall-tau correlation within Holmes (left) and compared to the OpenLLM benchmark (right). Green stars indicate significant correlations ($p<0.05$).
Figure 10: Effect of scaling LM parameters considering the T5 and Pythia model families providing eight and five different sizes. We address the overall scope (left) and the different types of linguistic phenomena (right).
iv) The effect of language model architecture.
Next, we discuss the impact of model architecture on the linguistic competence of LMs. In Figure 11 (left), we compare encoder and decoder LMs. Due to the absence of big encoder LMs, we consider five encoder and six decoder LMs with up to 220m parameters. Encoder LMs show a higher $mwr$ of $52\%$ than decoder LMs ($21\%$). This observation is the most saturated for morphology or syntax, encompassing a variety of token-level phenomena, like part-of-speech. We assume that the missing bi-directional encoding of decoder LMs causes this lower performance because the available context of one token heavily depends on its position. Thus, even common tokens, like the, have different potential representations - at the beginning or in the middle of a sentence. These instabilities are further evident when considering Figure 11 (right) which reports the accuracy for the top-20 most common POS tokens (such as the) based on the pos, xpos, upos dataset. Given their high frequency, one expects stable prediction performance. Surprisingly, encoder LMs (BERT and RoBERTa) show higher median accuracy and clearly lower deviations compared to the same-size decoder counterpart (GPT2). While scaling model size to 12B (Pythia) and 70B (Llama-2) allows for improved accuracy and lower deviations, decoder LMs do not match the encoder performance, even up to 700 times bigger.
Figure 11: Comparison of the phenomenon types for encoder and decoder LMs (left) and on the right, the accuracy of the top-20 most common tokens of the three part-of-speech probing datasets for BERT, RoBERTa, GPT2, Pythia, and Llama-2.
v) The effect of scaling parameters.
We discuss how the number of parameters influences the linguistic competence of LMs. Given the variety of LMs of different sizes, we focus on the Pythia (decoder-only) and T5 (encoder-decoder) families. From Figure 10, we observe for both Pythia and T5 that the linguistic competence scales with model size, and it is particularly pronounced after exceeding 0.5B (Pythia) and 1.0B (T5) parameters. Again, model architecture is crucial, as T5 LMs (encoder-decoder) exhibit a clearly higher mean winning rate of $40-70\%$ than Pythia (decoder-only) ones with $mwr$ of $20-60\%$. Further, we found formal phenomena evolving differently with increased model size than functional ones. Specifically, morphology and syntax start at a lower level, with an apparent performance jump after 0.5B (Pythia) and 1.0B (T5) parameters, followed by slow but steady growth. Differently, semantics, reasoning, and discourse start at a higher $mwr$, followed by a continuous improvement as the model size grows. From these results, we assume more parameters allow LMs to better approximate simpler co-occurrences in the near neighborhood of words to understand formal phenomena like word dependencies. In contrast, more parameters do not have the same pronounced effect on functional phenomena, like rhetorical relations, which require an LM to acquire more distant and complex word co-occurrences.
Model Morphology Syntax Semantics Reasoning Discourse Overall
Comparison against Llama-2 with 7 billion parameters
Llama-2-Chat -8% +3% -5% -9% -3% -2%
Comparison against T5 with 11 billion parameters
FLAN-T5 +10% +2% -2% +6% -2% +1%
Comparison against Pythia with 12 billion parameters
Dolly-v2 +4% -3% -9% -3% +4% -4%
Comparison against Llama-2 with 13 billion parameters
Tülu-2 +5% +2% -15% 0% -30% -8%
Orca-2 -1% -3% -4% +4% -5% -2%
Llama-2-chat +3% +1% -6% +3% -1% -1%
Vicuna-v1.5 +23% +7% -3% +6% -6% +4%
Comparison against UL2 with 20 billion parameters
FLAN-UL2 +40% +16% +7% +13% +1% +13%
Comparison against Mixtral with ~47 billion parameters
Mixtral-Instruct +4% +3% 0% +6% -2% +2%
Comparison against Llama-2 with 70 billion parameters
Tülu-2 +15% 0% -11% -3% 0% -2%
Llama-2-Chat +23% +14% +2% +4% +17% +10%
Average +10% +4% -3% +4% -2% +1%
Table 1: Effect of instruction tuning on the mean winning rate compared to the pre-trained LMs.
vi) The effect of instruction tuning.
Finally, we focus on how instruction tuning affects LMs’ linguistic competence and compare the tuned LMs with their base models—for example, FLAN-UL2 vs. UL2. From results in Table 1, we note less saturated effects for the overall scope while being more pronounced for the five phenomenon types - again emphasizing the structured and comprehensive evaluation of linguistic competence. On average, we found instruction tuning has the highest effect on morphology ($+10\%$) followed by syntax ($+4\%$), reasoning ($+4\%$), and a negative effect for semantics $-3\%$ and discourse $-2\%$. These results confirm previous assumptions that instruction tuning updates are often superficial (Yadav et al., 2023; Hershcovitch et al., 2024; Sharma et al., 2023) and that LMs are better at mimicking language (formal phenomena) than understanding it, measured with functional phenomena (Mahowald et al., 2024). Further, larger models benefit more from instruction tuning. Llama-2-70b-Chat and FLAN-UL2 gain up to $+23\%$ and $+40\%$ for morphology and $+10\%$ and $+13\%$ on average. In addition, decoder-only LMs (Llama-2 and Pythia) tend to show less pronounced positive effects than encoder-decoder LMs (FLAN-T5-XXL and FLAN-UL2). However, they better understand reasoning phenomena. When comparing LMs based on Llama-2-13b, we see that specific fine-tuning methods shape the LMs differently. The top-ranked 13b LM for Holmes and OpenLLM, Vicuna, was trained on 125k instructions, less than other models. Thus, high quality is more important than the number of instructions for LMs’ linguistic competence. Tülu loses performance while being trained on a large mixture of data (approx. 330k instructions), the same for its 70b version. Finally, the focus of Orca-2 on reasoning is also reflected in its embedding space. These insights show again that while providing a particular perspective, Holmes shows clear differences between LMs and allows us to map them to methodological decisions.
6 Efficiency
[…]
Besides excluding licensed data (18 probing datasets), we analyze the effect of discarding training instances. As a result, we reduce the computation for encoding and the actual probing simultaneously. We follow Perlitz et al. (2023) and calculate the rank resolution, $95\%$ CI of model rank difference. This measure indicates the maximum expected rank deviation from evaluating an LM on FlashHolmes compared to Holmes. For example, a rank resolution of one means that an LM evaluated on FlashHolmes and Holmes has the same rank or switch place with its neighbors with a probability of 95%. Figure 12 shows the resulting rank resolution when training only on a fraction of the instances, from $1/2$ to $1/512$. Solely focusing on efficiency ($1/512$) still provides a decent rank resolution of ~2.7. In contrast, considering $1/2$ of the training data results in the best reliability of ~1.0. To balance benchmark reliability and efficiency, we compose FlashHolmes using $1/32$ of the training instances. Precisely, it reduces the computation expenses of evaluating LMs to ~3% of what Holmes would have required while preserving a high rank-correlation of ~1.3.
Figure 12: Analysis of the reliability vs. efficiency trade-off when reducing the number of training data.
7 Related Work
Benchmarking LMs
Benchmarks approximate LMs abilities like general language understanding (Wang et al., 2019b, a), out-of-distribution generalization (Yang et al., 2023; Waldis et al., 2024b), adversarial scenarios (Nie et al., 2020; Wang et al., 2021), or retrieval like BEIR (Thakur et al., 2021) or MTEB (Muennighoff et al., 2023). With the advent of larger LMs, the methodological focus shifted to prompting-based evaluations which evaluate the LMs’ response to provided instructions (Brown et al., 2020; Hendrycks et al., 2021; Srivastava et al., 2022) covering application-oriented tasks (Liang et al., 2023), or mathematical reasoning (e.g., GSM8K Cobbe et al. (2021)).
Assessing the Linguistic Competence of LMs
The analysis of LMs’ linguistic competence ranges from analyzing static word vectors (Köhn, 2015), sentence embeddings (Conneau et al., 2018; Adi et al., 2017), the internals of translation models (Shi et al., 2016; Bau et al., 2019), or contextualized LMs (Tenney et al., 2019b, a; Hewitt and Manning, 2019). Other work addressed methodological aspects, such as using control tasks (Hewitt and Liang, 2019), assessing LMs from an information theory perspective (Voita and Titov, 2020; Pimentel et al., 2020), or evaluating causal effects in LMs (Elazar et al., 2021). Finally, another line of work focuses on whether LMs follow human understanding of linguistic competence when solving downstream tasks (Belinkov, 2022; Aw et al., 2023; Mahowald et al., 2024). However, Mosbach et al. (2020b) and Waldis et al. (2024a) found fine-tuning for downstream tasks actually hurting the understanding of linguistic phenomena.
While prior studies assessing the linguistic competence of LMs tend to focus on a limited set of linguistic phenomena or models, Holmes provides extensive coverage of both phenomena and evaluated LMs. Unlike recent evaluations based on prompting methods (Blevins et al., 2023; Liang et al., 2023; Amouyal et al., 2024), Holmes assesses the internal representations of LMs directly. This approach allows for detailed analysis of specific model characteristics, such as architecture, and helps separate the linguistic competence from other cognitive abilities. Thereby, we respond to recent calls for a thorough and explicit evaluation of linguistic phenomena (Hu and Levy, 2023; Lu et al., 2023; Mahowald et al., 2024).
8 Conclusion
Holmes marks the most up-to-date and extensive consolidation of existing resources addressing the need to assess the linguistic competence of LMs in isolation. Our experiments demonstrate that LMs’ linguistic competence is pronounced regarding formal phenomena but lacks functional ones when information about broader textual contexts, such as rhetorical structure, is required. Further, size, architecture, and instruction tuning crucially account for differences among LMs. As LM and resources in the landscape of linguistics continue to grow, we will actively extend Holmes with further probing datasets, evaluate upcoming LMs, and plan to incorporate multilingualism.
Ethical Considerations and Limitations
Language
Holmes as well as FlashHolmes solely assess linguistic phenomena for the English language. As we plan to expand the benchmark and scope of multilingual data, we focus momentarily on English because of the widespread availability of resources, including curated corpora and the diversity of available LMs.
Phenomena and LM Coverage
We agree with Liang et al. (2023) and see one fundamental aspect in composing a benchmark in acknowledging its incompleteness. Both linguistic phenomena and LMs are a subset of the variety of available resources. We consolidated them carefully to provide a comprehensive scope of the linguistic competence and various LMs. However, as benchmarks evolve as tools to assess LMs, we will further expand Holmes both with the existing and upcoming LMs and data resources.
Data Availability
Linguistic annotations, in particular more complex ones targeting phenomena like discourse, are money and time-wise expensive. Out of 208 datasets included in Holmes, 18 probing datasets are based on licensed resources and are not freely available. However, with FlashHolmes, we provide an effective and efficient alternative based on open-access resources. Furthermore, upon confirming the granted access, we are happy to share our probing datasets, including those based on the licensed resources.
Bias
As Holmes relies on existing resources, it inherits the bias embodied in this data. Examples of such bias are gender equality or gender fairness, like the use of neo pronouns such as em in Lauscher et al. (2023).
Appendix A Additional Details of Holmes
A.1 Additional Details on the Evolution of Probing Literature
We analyze publication trends by year and venue as shown in Table 2. Less work was published between 2015-2018 (earlier) focusing on LSTM-based (Linzen et al., 2016; Conneau et al., 2018) and static LMs (Köhn, 2015; Linzen et al., 2016; Belinkov et al., 2017; Conneau et al., 2018). With the release of BERT (Devlin et al., 2019) in 2019, we note increasing attention to analyzing linguistic abilities within LMs, with a peak of 90 papers in 2022.444Note that EMNLP-23 and AACL-23 proceedings were not published when conducting this meta-analysis. Considering the venue, more than half of the relevant work (149 papers) was published at major conferences (ACL and EMNLP), and 68 papers were published at AACL, EACL, NAACL, and COLING. [^5] Note that EMNLP-23 and AACL-23 proceedings were not published when conducting this meta-study. In addition, we observe a constant contribution of TACL, various workshops, such as Analyzing and Interpreting Neural Networks for NLP or Representation Learning for NLP.
A.2 Experimental Details
Probing Hyperparameters
Following previous work (Hewitt and Liang, 2019; Voita and Titov, 2020), we use fixed hyperparameters for training the probes: 20 epochs, where we find the best one using dev instances; AdamW (Loshchilov and Hutter, 2019) as optimizer; a batch size of 64; a learning rate of 0.0005; a dropout rate of 0.2; a warmup rate of 10% of the steps; random seeds: $[0,1,2,3,4]$
Hardware
We run all of our experiments using 12 Nvidia RTX A6000 GPUs. Every GPU provides 48GB of memory and 10752 CUDA Cores.
Considered LMs
Table 8 outlines the details of the LMs we evaluate on Holmes in this work.
A.3 Linguistic Task Categorization
We show in Table 3, Table 4, Table 7, Table 5, and Table 6 which resources Holmes use to cover morphology, syntax, semantics, reasoning, and discourse phenomena. This includes 33 works providing the data, the specific linguistic phenomena, or both. For example, for readability we use the data of Weischedel et al. (2013) and calculated the flesch score (Flesch, 1948).
Morphology
First, we feature 19 tasks verifying morphology phenomena: Anaphor agreement, determiner noun agreement, subject-verb agreement and irregular forms (Warstadt et al., 2020; Huebner et al., 2021).
Syntax
The second group of 75 tasks verifies the following syntax phenomena: Part-of-speech tagging and constituent labeling (Weischedel et al., 2013); dependency labeling (Silveira et al., 2014); bigram-shift, tree-depth, top-constituent-task, and sentence-length (Conneau et al., 2018); subject- & object-number, and deoncausative-inchoative alternation based on Klafka and Ettinger (2020); binding, control/raising, negative polarity item licensing, island-effects, argument-structure, ellipsis, and filler-gap (Warstadt et al., 2020; Huebner et al., 2021).
earlier 2019 2020 2021 2022 2023 Total
ACL 2 10 12 9 34 25 92
AACL - - - - 1 - 1
COLING - - 10 - 9 - 19
EACL - - - 7 - 15 22
EMNLP 2 4 13 17 21 - 57
NAACL - 3 - 9 14 - 26
TACL 1 1 2 3 3 1 11
Workshops 4 4 10 10 7 1 36
Other 1 2 1 1 1 4 10
Probing 10 24 48 56 90 46 274
All Papers 8,056 3,111 3,822 4,294 5,133 3,647 28,063
Table 2: Evolution of probing studies. Note that EMNLP-23 and AACL-23 proceedings were not published when conducting this meta-study.
Semantics
Third, consider 67 tasks covering semantics phenomena: Named-entity labeling and semantic-role labeling (Weischedel et al., 2013); subject- and object-number, tense, semantic odd man out, word content, and coordination inversion (Conneau et al., 2018); semantic relation classification (Hendrickx et al., 2010); semantic proto-roles (Rudinger et al., 2018a); factuality (Rudinger et al., 2018b); genericity (Govindarajan et al., 2019); event structure (Gantt et al., 2022); time (Vashishtha et al., 2019); word sense (White et al., 2016); sentiment analysis (Socher et al., 2013); object- and subject-animacy, object- and subject-gender, verb-tense, and verb-dynamic Klafka and Ettinger (2020); metaphor (Mohler et al., 2016; Birke and Sarkar, 2006; Steen et al., 2010); complex word identification (Paetzold and Specia, 2016); and passive (Krasnowska-Kieraś and Wróblewska, 2019). In addition, we derive synonym-/antonym-detection task using WordNet (Miller, 1995) and the texts from OntoNotes v5 Weischedel et al. (2013).
Reasoning
Forth, 19 tasks cover reasoning phenomena: Paraphrasticity with negation and antonyms (Vahtola et al., 2022); negation detection (Szarvas et al., 2008; Konstantinova et al., 2012; Morante and Blanco, 2012); negation-span classification (Szarvas et al., 2008; Konstantinova et al., 2012); negation-correspondence (Szarvas et al., 2008; Konstantinova et al., 2012); speculation detection, speculation-span classification, and speculation-correspondence (Szarvas et al., 2008); and always-never, age comparison, objects comparison, antonym negation, property conjunction, taxonomy connection, encyclopedic composition, and multi-hop composition (Talmor et al., 2020).
Discourse
Finally, Holmes embodies 28 task addressing discourse phenomena: Co-reference resolution Weischedel et al. (2013); bridging (Hou, 2018, 2020; Pandit and Hou, 2021); discourse connective (Nie et al., 2019); sentence order and next-sentence prediction (Narayan et al., 2018); discourse correspondence, discourse order, discourse relation, discourse distance, discourse explicit classes, discourse implicit classes (Webber et al., 2019; Kurfalı and Östling, 2021); and rst-count/-depth/-distance/-relation/-relation-group/-successively/-type (Carlson et al., 2001; Koto et al., 2021; Kurfalı and Östling, 2021; Zeldes, 2017).
A.4 Details of Probing Dataset Composition
Whenever possible, we rely on established probing datasets and transform instances into a unified format: 1) an input $x$ which is either one or a pair of span(s) or sentence(s), including the string and an optional starting and ending index in the context $c$ when task type is either a span or span-pair classification; 2) an optional textual context $c$ to encode $x$, for example the sentence in which a span occurs; and 3) a corresponding label $y$. If given, we use the original train/dev/test splits. However, if this division does not exist, we use a 70/10/20 ratio to form these splits. Furthermore, we adapt the design of some tasks to map to our task format. Exemplary, for the oLMmpics (Talmor et al., 2020) dataset, we transform the mask-filling tasks into a binary classification where the correct label corresponds to a sentence with a correctly filled mask and incorrect to a sentence where the mask was filled wrongly.
OnToNotes
Following Tenney et al. (2019b, a), we use the OntoNotes (Weischedel et al., 2013) dataset to derive part-of-speech tagging, constituent labeling, named-entity labeling, semantic role, and co-reference resolution probing datasets. Further, we consider with constituent maximum depth and constituent node length further properties of the constituent tree this dataset OntoNotes.
Dependency Corpus
As in Tenney et al. (2019b, a), we use Universal Dependencies annotations of the English Web Treebank to form a dependency labeling datasets.
Context Probes
Presented in Klafka and Ettinger (2020), we compose nine datasets to verify information about context words.
BLiMP Dataset
Using the data presented in the BLiMP benchmark (Warstadt et al., 2020), we derive 67 probing datasets verifying specific phenomena, like island effect, covering morphology, syntax, and semantics. Unlike the original version, we compose a binary classification task for every phenomenon. Precisely, whether to accept or reject a given sentence, where rejecting means that the given linguistic phenomena is violated.
Zorro Dataset
As for the BLiMP tasks, we convert the 21 distinct Zorro tasks into a binary classification task on whether a sentence accepts or rejects the given linguistic phenomena is violated.
SemEval-2010 Task 8
For semantic relation classification we rely on the dataset of Hendrickx et al. (2010).
Decompositional Semantics Initiative
The Decompositional Semantics Initiative [^6] https://decomp.io/ provides a large number of datasets to verify semantic phenomena. Apart of the common use semantic proto-roles (Rudinger et al., 2018a), we use their collection of works to compose probing datasets for factuality (Rudinger et al., 2018b), genericity (Govindarajan et al., 2019), event structure (Vashishtha et al., 2019), time (Vashishtha et al., 2019), and word sense (White et al., 2016).
Sentiment Analysis
We use the commonly used work of Socher et al. (2013) and form a probing dataset targeting sentiment.
Metaphor
As in Aghazadeh et al. (2022), we use the data from Mohler et al. (2016); Birke and Sarkar (2006); Steen et al. (2010) to form three metaphor datasets.
Complex Word Identification
We consider word complexity for the first time and use the data presented in Paetzold and Specia (2016). It provides annotations for different complexity levels of words.
Passive
We use data from Krasnowska-Kieraś and Wróblewska (2019) to form a probing dataset assessing knowledge about passive language.
Synonym / Antonym Replacement
Using the text of the OntoNotes (Weischedel et al., 2013) and Wordnet (Miller, 1995), we form a probing dataset to detect synonym and antonym replacement. Specifically, the binary classification task is: given two texts (the original and an updated one), was the updated one changed by replacing a word with its synonym or antonym?
Negation
With this work, we verify for the first time negation based on human annotated datasets (Vahtola et al., 2022; Szarvas et al., 2008; Konstantinova et al., 2012). Specifically, we form different probing datasets.
  • Is a text negated or not?
  • Given two text spans, does the negation within the first one correspond to the second one?
  • Given a text span, is it the cue or the scope of the negation?
oLMmpics
We form probing datasets addressing different lexical reasoning using the data presented in Talmor et al. (2020). As they provide multiple choices, we form correct instances by filling the gap with the correct option and wrong ones by filling in the other options. Specifically, we form dataset for always-never, age comparison, objects comparison, antonym-negation, multi-hop composition property conjunction, taxonomy conjunction, and encyclopedic composition.
Bridging
We rely on the data presented in Pandit and Hou (2021) and form two probing datasets. One is to verify whether a text is linguistically applicable, considering bridging (antecedent matches anaphora). And a second one to verify whether an antecedent and anaphora match.
Discourse Connective
Using data from Nie et al. (2019), we form a probing dataset to assess whether a given connective marker matches the discourse of the given text.
Sentence Order and Next Sentence Prediction
Following Narayan et al. (2018), we form two datasets to verify the order of good or badness of a given sentence and whether two sentences occur after each other.
Discourse Representation Theory
We use data from Webber et al. (2019) to compose eight probing datasets addressing discourse representation theory:
  • Four probing dataset predicting the class of a given span. We distinguish between implicit, explicit, implicit-coarse, and explicit-coarse.
  • The absolute distance, number of words, between two spans in the text.
  • Whether the order of two spans is correct or not.
  • Whether two spans have discourse relation or not.
  • The specific discourse relation of two spans.
Rhetorical Structure Theory
Using annotations from Carlson et al. (2001); Zeldes (2017), we compose 14 probing datasets addressing rhetorical theory. Specifically, we compose the following seven types of datasets for both works:
  • The rhetorical type of a text span, either nucleus or satellite.
  • The number of children of a text span within the rhetorical tree of the text.
  • The depth of a text span within the rhetorical tree of the text.
  • The number of edges between two text spans within the rhetorical tree.
  • The specific rhetorical relation between two text spans like conclusion.
  • The relation group of a specific rhetorical relation between two text spans like evaluation for the relation conclusion.
  • Whether two text spans occur after each other in the rhetorical tree.
Phenomena
Text
Text-Pair
Span
Span-Pair
Warstadt et al. (2020)
Huebner et al. (2021)
anaphor agreement 3
determiner noun agreement 10
irregular forms 3
subject-verb agreement 10
Table 3: Overview of resources and linguistic phenomena mapping for morphology. It shows the number of datasets for the phenomena by dataset type.
Phenomena
Text
Text-Pair
Span
Span-Pair
Weischedel et al. (2013)
Silveira et al. (2014)
Conneau et al. (2018)
Flesch (1948)
Klafka and Ettinger (2020)
Warstadt et al. (2020)
Huebner et al. (2021)
argument-structure 20
bigram-shift 1
binding 8
case 1
constituent parsing 2 1
control/raising 5
deoncausative-inchoative alternation 1
dependency parsing 1
ellipsis 3
filler-gap 9
island-effects 10
local attractor 1
object-number 2
part-of-speech 3
readability 1
sentence-length 1
subject-number 2
top-constituent-task 1
tree-depth 1
Table 4: Overview of resources and linguistic phenomena mapping for syntax. It shows the number of datasets for the phenomena by dataset type.
Phenomena
Text
Text-Pair
Span
Span-Pair
Vahtola et al. (2022)
Szarvas et al. (2008)
Konstantinova et al. (2012)
Morante and Blanco (2012)
Talmor et al. (2020)
age comparison 1
always-never 1
antonym negation 1
encyclopedic composition 1
multi-hop composition 1
negation 3 1 2 2
objects comparison 1
property conjunction 1
speculation 1 1 1
taxonomy connection 1
Table 5: Overview of resources and linguistic phenomena mapping for reasoning. It shows the number of datasets for the phenomena by dataset type.
Phenomena
Text
Text-Pair
Span
Span-Pair
Weischedel et al. (2013)
Pandit and Hou (2021)
Nie et al. (2019)
Narayan et al. (2018)
Webber et al. (2019)
Carlson et al. (2001)
Zeldes (2017)
bridging 1 1
co-reference resolution 1
discourse connective 1
discourse representation theory 8
next-sentence prediction 1
rethorical structure theory 6 8
sentence order 1
Table 6: Overview of resources and linguistic phenomena mapping for discourse. It shows the number of datasets for the phenomena by dataset type.
Phenomena
Text
Text-Pair
Span
Span-Pair
Weischedel et al. (2013)
Conneau et al. (2018)
Klafka and Ettinger (2020)
Warstadt et al. (2020)
Huebner et al. (2021)
Hendrickx et al. (2010)
Rudinger et al. (2018a)
Rudinger et al. (2018b)
Govindarajan et al. (2019)
Gantt et al. (2022)
Vashishtha et al. (2019)
White et al. (2016)
Socher et al. (2013)
Mohler et al. (2016)
Birke and Sarkar (2006)
Steen et al. (2010)
Paetzold and Specia (2016)
Krasnowska-Kieraś and Wróblewska (2019)
Miller (1995)
complex word identification 1
coordination inversion 1
event structure 4 2
factuality 1
genericity 6
metaphor 4
named-entity labeling 1
negative polarity item licensing 4
object-animacy 1
object-gender 1
passive 1
quantifiers 6
semantic relation classification 1
semantic proto-roles 20
semantic odd man out 1
semantic-role labeling 1
sentiment analysis 1
subject-animacy 1
subject-gender 1
synonym-/antonym-detection 1
tense 2
time 1
verb-dynamic 1
word content 1
word sense 1
Table 7: Overview of resources and linguistic phenomena mapping for semantics. It shows the number of datasets for the phenomena by dataset type.
Model Citation Size Pre-Training Objective Pre-Training Data Huggingface Tag
Encoder-Only Language Models
ALBERT Lan et al. (2020) 10 million MLM+SOP 16GB albert-base-v2
BERT Tenney et al. (2019a) 110 million MLM+NSP 16GB bert-base-uncased
DeBERTa He et al. (2021) 100 million MLM 80GB microsoft/deberta-base
DeBERTa-v3 He et al. (2023) 86 million MLM+DISC 160GB microsoft/deberta-v3-base
ELECTRA Clark et al. (2020) 110 million MLM 16GB google/electra-base-discriminator
RoBERTa Liu et al. (2019) 110 million MLM+DISC 160GB roberta-base
Decoder-Only Language Models
GPT2 Radford et al. (2019) 117 million LM 40GB gpt2
Pythia-70m Biderman et al. (2023) 70 million LM 300 billion tokens EleutherAI/pythia-70m
Pythia-160m Biderman et al. (2023) 160 million LM 300 billion tokens EleutherAI/pythia-160m
Pythia-410m Biderman et al. (2023) 410 million LM 300 billion tokens EleutherAI/pythia-410m
Pythia-1b Biderman et al. (2023) 1 billion LM 300 billion tokens EleutherAI/pythia-1B
Pythia-1.4b Biderman et al. (2023) 1.4 billion LM 300 billion tokens EleutherAI/pythia-1.4B
Pythia-2.8b Biderman et al. (2023) 2.8 billion LM 300 billion tokens EleutherAI/pythia-2.8B
Pythia-6.9b Biderman et al. (2023) 6.9 billion LM 300 billion tokens EleutherAI/pythia-6.9B
Pythia-12b Biderman et al. (2023) 12 billion LM 300 billion tokens EleutherAI/pythia-12B
Pythia-70m-dedup Biderman et al. (2023) 70 million LM 207 billion tokens EleutherAI/pythia-70m-deduped
Pythia-160m-dedup Biderman et al. (2023) 160 million LM 207 billion tokens EleutherAI/pythia-160m-deduped
Pythia-410m-dedup Biderman et al. (2023) 410 million LM 207 billion tokens EleutherAI/pythia-410m-deduped
Pythia-1b-dedup Biderman et al. (2023) 1 billion LM 207 billion tokens EleutherAI/pythia-1B-deduped
Pythia-1.4b-dedup Biderman et al. (2023) 1.4 billion LM 207 billion tokens EleutherAI/pythia-1.4B-deduped
Pythia-2.8b-dedup Biderman et al. (2023) 2.8 billion LM 207 billion tokens EleutherAI/pythia-2.8B-deduped
Pythia-6.9b-dedup Biderman et al. (2023) 6.9 billion LM 207 billion tokens EleutherAI/pythia-6.9B-deduped
Pythia-12b-dedup Biderman et al. (2023) 12 billion LM 207 billion tokens EleutherAI/pythia-12B-deduped
Dolly-v2 Conover et al. (2023) 12 billion LM+IT 300 billion token + 15K instructions databricks/dolly-v2-12b
Llama-2-7b Touvron et al. (2023) 7 billion LM 2.4 trillion tokens meta-llama/Llama-2-7b-hf
Llama-2-13b Touvron et al. (2023) 13 billion LM 2.4 trillion tokens meta-llama/Llama-2-13b-hf
Llama-2-70b Touvron et al. (2023) 70 billion LM 2.4 trillion tokens meta-llama/Llama-2-70b-hf
Llama-2-7b-chat Touvron et al. (2023) 7 billion LM+IT 2.4 trillion tokens + 27,5K instructions meta-llama/Llama-2-7b-chat-hf
Llama-2-13b-chat Touvron et al. (2023) 13 billion LM+IT 2.4 trillion tokens + 27,5K instructions meta-llama/Llama-2-13b-chat-hf
Llama-2-70b-chat Touvron et al. (2023) 70 billion LM+IT 2.4 trillion tokens + 27,5K instructions meta-llama/Llama-2-70b-chat-hf
IBM-Merlinite Sudalairaj et al. (2024) 7 billion LM+IT 2.4 trillion tokens + 1400k instructions ibm/merlinite-7b
IBM-Labradorite Sudalairaj et al. (2024) 13 billion LM+IT 2.4 trillion tokens + 1400k instructions ibm/labradorite-13b
Vicuna-13b-v1.5 Zheng et al. (2023) 13 billion LM+IT 2.4 trillion tokens + 125k instructions lmsys/vicuna-13b-v1.5
Orca-2-13b Mitra et al. (2023) 13 billion LM+IT 2.4 trillion tokens + 817K instructions microsoft/Orca-2-13b
Wizard-13B-v1.2 Xu et al. (2023) 13 billion LM unknown WizardLM/WizardLM-13B-V1.2
Tülu-2-13b Wang et al. (2023) 13 billion LM+IT 2.4 trillion tokens + 330k instructions allenai/tulu-2-13b
Tülu-2-dpo-13b Wang et al. (2023) 13 billion LM+IT 2.4 trillion tokens + 330k instructions tulu-2-dpo-13b
Tülu-2-70b Wang et al. (2023) 70 billion LM+IT 2.4 trillion tokens + 330k instructions allenai/tulu-2-70b
Tülu-2-dpo-70b Wang et al. (2023) 70 billion LM+IT 2.4 trillion tokens + 330k instructions tulu-2-dpo-70b
Mistral-7b Jiang et al. (2023) 7 billion LM unknown mistralai/Mistral-7B-v0.1
Mistral-7b-Inst Jiang et al. (2023) 7 billion LM unknown mistralai/Mistral-7B-Instruct-v0.1
Mixtral-8x7b Jiang et al. (2024) 47 billion LM unknown mistralai/Mixtral-8x7B-v0.1
Mixtral-8x7b-Inst Jiang et al. (2024) 47 billion LM unknown mistralai/Mistral-7B-v0.1
Encoder-Decoder Language Models
BART Lewis et al. (2020) 121 million DAE 160GB google/facebook/bart-base
T5-small Raffel et al. (2020) 60 million DAE 800GB google/t5-small-lm-adapt
T5-base Raffel et al. (2020) 220 million DAE 800GB google/t5-base-lm-adapt
T5-large Raffel et al. (2020) 770 million DAE 800GB google/t5-large-lm-adapt
T5-xl Raffel et al. (2020) 3 billion DAE 800GB google/t5-xl-lm-adapt
T5-xxl Raffel et al. (2020) 11 billion DAE 800GB google/t5-xxl-lm-adapt
FLAN-T5-small Raffel et al. (2020) 60 million DAE+IT 800GB + 1.8k tasks google/t5-small-lm-adapt
FLAN-T5-base Raffel et al. (2020) 220 million DAE+IT 800GB + 1.8k tasks google/t5-base-lm-adapt
FLAN-T5-large Raffel et al. (2020) 770 million DAE+IT 800GB + 1.8k tasks google/t5-large-lm-adapt
FLAN-T5-xl Raffel et al. (2020) 3 billion DAE+IT 800GB + 1.8k tasks google/t5-xl-lm-adapt
FLAN-T5-xxl Raffel et al. (2020) 11 billion DAE+IT 800GB + 1.8k tasks google/t5-xxl-lm-adapt
TK-Instruct Wang et al. (2022) 11 billion billion DAE+IT 800GB + 1.6k tasks allenai/tk-instruct-11b-def
UL2 Tay et al. (2023) 20 billion DAE 800GB google/ul2
FLAN-UL2 Tay et al. (2023) 20 billion DAE+IT 800GB + 100k instructions google/flan-ul2
Static Language Models
Glove-6B Pennington et al. (2014) - WP 6 billion tokens glove.6B.300d
Glove-840B Pennington et al. (2014) - WP 840 billion tokens glove.840B.300d
Table 8: Overview of the evaluated LMS covering the corresponding citation, model size, model architecture, pre-training objective & data, and the Huggingface model tag. Regarding the pre-training objective, we distinguish between masked language modeling (MLM), sentence order prediction (SOP), next sentence prediction (NSP), next word prediction (LM), instruction fine-tuning (IT), word denoising (DAE), and word probabilities from word co-occurrences (WP). For pre-training data, we report known numbers, either as the size of the corpora in gigabytes (GB), the number of pre-training tokens, the number of instructions for fine-tuning, or the number of tasks for instruction fine-tuning.